# SPACE: SPike-Aware Consistency Enhancement for Test-Time Adaptation in Spiking Neural Networks

### v1

@@ 6 Kernel-Based Consistency Regularization
Shifted Domain w.o. TTA SPACE SPACE + Kernel Embedding
noise gaussian 72.38% 77.98% 77.85%
shot 74.70% 79.34% 79.34%
speckle 71.75% 77.38% 77.33%
impulse 58.57% 69.41% 69.25%
blur defocus 63.05% 71.59% 71.42%
gaussian 57.03% 68.31% 67.98%
motion 64.44% 72.14% 72.21%
zoom 71.33% 74.67% 74.61%
weather snow 76.32% 78.43% 78.48%
fog 43.57% 52.80% 52.46%
frost 75.72% 79.59% 79.66%
digital brightness 82.44% 83.22% 83.25%
contrast 22.54% 23.85% 23.40%
elastic_transform 75.01% 75.49% 75.42%
pixelate 70.25% 76.24% 76.13%
jpeg_compression 84.28% 82.88% 82.98%
spatter 78.42% 77.52% 77.47%
saturate 80.85% 82.54% 82.64%
Average 67.93% 72.41% 72.41%
Table 5: Performance comparison of SPACE without and with kernel embedding on CIFAR-10-C dataset at level 5 corruption, using the VGG9 architecture.
Shifted Domain w.o. TTA MEMO SPACE
DropPixel 44.53% 43.94% 46.97%
DropEvent 50.00% 51.52% 53.41%
RefractoryPeriod 35.16% 34.85% 39.39%
TimeJitter 42.19% 40.91% 42.42%
SpatialJitter 83.98% 82.95% 85.23%
UniformNoise 75.39% 74.24% 75.39%
Average 55.21% 54.74% 57.14%
Table 6: Performance comparison of different methods on DVS Gesture-C dataset at level 5 corruption.
@@
To evaluate the impact of incorporating kernel embedding into the loss function, we tested its performance on CIFAR-10-C dataset, using the BNTT model with the VGG9 architecture, as shown in Tab. 5. The results indicate that adding kernel embedding provides little to no improvement over the original SPACE method, with the average performance remaining nearly identical. This limited effectiveness can be attributed to several factors. First, the distribution differences introduced by augmentations, such as noise, blur, and weather effects, are relatively small, reducing the need for advanced alignment mechanisms like kernel embedding. Additionally, the existing similarity measures in SPACE already effectively capture the relationships between augmented samples, leaving little room for further optimization. Moreover, the added computational complexity of kernel embedding, especially the cost of kernel calculations, may even slightly hinder performance in scenarios where the original approach is already sufficient. These factors combined suggest that kernel embedding is not particularly beneficial in this experimental setting.
7 Performance on DVS Gesture-C Dataset
SNNs are inherently suited for processing event-based neuromorphic data due to their temporal dynamics and energy efficiency. Static image datasets like CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C are commonly used to evaluate robustness to corruption and distribution shifts, but they lack the temporal and asynchronous characteristics of real-world event-based data. DVS-Gesture [1] allows us to evaluate the proposed method’s performance on neuromorphic datasets, emphasizing its compatibility with dynamic vision tasks and data with temporal structure.
@@ 7 Performance on DVS Gesture-C Dataset
Following the state-of-the-art methodology [18], we evaluate the robustness improvements achieved by SPACE using DVS-Gesture-C, a corrupted variant of the standard DVS-Gesture dataset. This variant introduces six distinct corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise. These corruptions, implemented via the Tonic API [28], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies. To ensure a comprehensive and stringent assessment, we consistently employ the highest severity level across all experimental evaluations. The pre-trained model used in this study follows a 2CONV-2FC architecture, which was selected for its balance between computational efficiency and representational capacity, ensuring a fair and reliable comparison.
The experimental results demonstrate that the proposed SPACE method achieves superior performance on the DVS-Gesture-C dataset compared to both w.o. TTA and the MEMO method. As shown in Tab. 6, SPACE consistently outperforms w.o. TTA across all six corruption types, with the largest improvements observed for RefractoryPeriod (+4.23%) and DropEvent (+3.41%). Compared to MEMO, SPACE also achieves higher accuracy in most scenarios, with notable gains for RefractoryPeriod (+4.54%) and DropPixel (+3.03%). On average, SPACE achieves an accuracy of 57.14%, significantly surpassing both w.o. TTA (55.21%) and MEMO (54.74%).
While MEMO improves TTA performance by enforcing consistency among augmented samples, it faces significant challenges in SNNs due to the loss of temporal information when relying on backpropagation. This limitation is evident in the results, where MEMO struggles to handle certain corruption types effectively, such as RefractoryPeriod and TimeJitter. In contrast, SPACE addresses these challenges by fully leveraging the temporal dynamics and event-based nature of SNNs, resulting in more robust and reliable performance across diverse corruption types. These results highlight the robustness and effectiveness of SPACE in improving model performance under challenging conditions, making it a highly suitable method for event-based data with various noise and corruption types.

### v2

@@ Appendix A Kernel-Based Consistency Regularization
To evaluate the impact of incorporating kernel embedding into the loss function, we tested its performance on CIFAR-10-C dataset, using SNN-VGG9 model, as shown in Table 3. The results indicate that adding kernel embedding provides little to no improvement over the original SPACE method, with the average performance remaining nearly identical. This limited effectiveness can be attributed to several factors. First, the distribution differences introduced by augmentations, such as noise, blur, and weather effects, are relatively small, reducing the need for advanced alignment mechanisms like kernel embedding. Additionally, the existing similarity measures in SPACE already effectively capture the relationships between augmented samples, leaving little room for further optimization. Moreover, the added computational complexity of kernel embedding, especially the cost of kernel calculations, may even slightly hinder performance in scenarios where the original approach is already sufficient. These factors combined suggest that kernel embedding is not particularly beneficial in this experimental setting.
Table 3: Performance comparison of SPACE without/with kernel embedding (k.e.) on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%). The bold number indicates the best result.
Method Noise Blur Weather Digital Average
Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
SPACE + k.e. 77.85 79.34 69.25 71.42 67.64 72.21 74.61 78.48 52.46 79.66 83.25 23.40 75.42 76.13 82.98 70.94
Appendix B Exploration of Alternative Alignment Objectives
B.1 Two Alternative Alignment Objectives
Average Membrane Potential (SPACE + a.m.p.)
In the main paper, we use feature maps based on the spike counts over the entire time window as the alignment objective. In this section, we explore an alternative biologically inspired metric: the average membrane potential obtained from the LIF neurons in the final layer of the feature extractor. Specifically, for a given test point $\mathbf{x}_{i}$, the membrane potentials are captured in the feature map $\mathbf{U}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times D}$, where $T$, $C$, and $D$ represent the time steps, the number of channels, and the spatial dimensions, respectively. For LIF neurons that spike at a specific time step, we use the membrane potential after the LIF reset, which indirectly captures the spiking information. To obtain the primary objective feature map $\mathbf{F}(\mathbf{x}_{i})$, we compute the temporal average of $\mathbf{U}(\mathbf{x}_{i})$ along the first dimension. The resulting feature map is then used to calculate the loss following Equations 4 and Equation 5 in the main paper. Finally, the model is updated via SGD.
Spikes through All Time Steps (SPACE + s.t.t.)
Instead of summing the spike counts over the temporal dimension, we explicitly retain the temporal information, resulting in a feature map with an additional dimension that captures the spiking activity at each time step. This approach preserves the precise temporal dynamics of the spikes and leverages them directly for alignment. Specifically, we align the channel-wise spiking activity across all time steps between each pair of augmented samples. The update rules remain consistent with those described in the main paper.
B.2 Experimental Results and Analysis
Table 4: Performance comparison of SPACE with different objectives on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%). The bold number indicates the best result. Average membrane potential and spike through all time steps are referred as a.m.p. and s.t.t. respectively.
Method Noise Blur Weather Digital Average
Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
SPACE + a.m.p. 74.87 76.05 61.64 64.90 65.79 65.88 72.22 77.66 45.66 76.64 82.81 21.57 75.53 70.91 84.74 67.79
SPACE + s.t.t. 72.74 73.84 57.86 62.91 63.66 64.33 70.94 76.35 43.16 74.77 82.30 21.20 74.24 69.35 83.97 66.11
We implement the two approaches above and evaluate on CIFAR-10-C dataset with SNN-VGG9 model. The results are shown in Table 4.
Compared to No Adapt, SPACE + a.m.p. improves the generalization ability of the model. However, its performance is noticeably inferior to SPACE on the majority of corruptions. This may be due to the fact that the membrane potentials of the same sample under different augmentations tend to exhibit similar patterns, making them less effective in capturing the temporal distribution characteristics of SNNs. For example, in extreme cases where the membrane potential of a LIF neuron stays near the firing threshold throughout all time steps without crossing it, SPACE + a.m.p. fails to capture this information. In contrast, SPACE can reflect this behavior through spike counts, which more directly encode spiking activity.
SPACE + s.t.t. fails to improve the generalization ability of the model. We hypothesize that this is due to the model overfitting during the alignment process. Specifically, the updates made to minimize the alignment loss may inadvertently lead to excessive specialization to the augmented samples, thereby hindering its ability to generalize effectively to unseen corruptions.
Appendix C Additional Ablation Study
Effect of Augmentation Quantity per Test Point
In the main paper, we follow MEMO [47] to determine the number of augmentations $M$. To analyze the effect of $M$ on our method, we evaluated the average accuracy and the evaluation time per test point the same computing resource conditions on CIFAR-10-C using SNN-VGG9 with varying augmentation quantities $M=\{2,4,8,16,32,64,128\}$. As shown in Table 5, at least 8 augmentations are required to achieve noticeable performance improvement. For $M=\{32,64,128\}$, while the performance continues to improve gradually, the evaluation time increases significantly. Considering the trade-off between accuracy and evaluation time, $M=32$ is the optimal choice.
Table 5: Performance comparison of SPACE with different augmentation quantities on CIFAR-10-C with SNN-VGG9 model regarding Accuracy (%) and Evaluation Time (seconds) per test point.
No Adapt Number of Augmentation
2 4 8 16 32 64 128
Accuracy 66.57 54.60 66.91 69.77 70.60 71.03 71.07 71.11
Evaluation Time 0.144 0.318 0.320 0.327 0.339 0.345 0.443 0.788
Appendix D Additional Experimental Information
D.1 More Details on Datasets
In this paper, we conducted six static datasets and one neuromorphic dataset to evaluate the out-of-distribution generalization ability. They are CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [15], ImageNet-V2 [35], ImageNet-R [17], ImageNet-A [18] and DVS Gesture-C [20].
CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C
These three datasets are commonly used benchmarks for evaluating the robustness of models under various types of input corruptions. They are derived from the original CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets by applying 15 corruption types (e.g., noise, blur, weather, and digital distortions) at 5 severity levels to the original CIFAR-10, CIFAR-100 [26], and Tiny-ImageNet [6] datasets, respectively. CIFAR-10-C and CIFAR-100-C consist of 10 and 100 classes, each with 50,000 training images and 10,000 testing images. Tiny-ImageNet-C, derived from Tiny-ImageNet, includes 200 classes with 100,000 training images and 10,000 validation images.
ImageNet V2/R/A
ImageNet-V2, ImageNet-R, and ImageNet-A are three datasets designed to evaluate model robustness under different real-world challenges. ImageNet-V2 consists of a new set of images collected under similar conditions as the original ImageNet [6], serving as a test of distribution shift. ImageNet-R focuses on image classification robustness across various artistic renditions, such as paintings, sketches, and cartoons, covering 200 ImageNet classes. ImageNet-A, on the other hand, is an adversarially filtered subset of natural images that models often misclassify, targeting their inherent weaknesses. Compared to corruption-based datasets, these benchmarks provide a better reflection of real-world scenarios and challenges.
DVS Gesture-C
@@ D.1 More Details on Datasets
DVS Gesture-C is a corrupted variant of the standard DVS Gesture [1] dataset. This variant introduces six distinct corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise. These corruptions, implemented via the Tonic API [29], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies.
D.2 More Details on Pretrained Models
SNN-VGG with BNTT
We adopt SNN-VGG with Batch-Normalization Through Time (BNTT) [24] as our main backbone model. BNTT decouples the parameters along the time axis within each layer, effectively capturing the temporal dynamics of spikes. The temporally evolving learnable parameters in BNTT enable neurons to regulate their spike rates across different time steps, facilitating low-latency and low-energy training from scratch. The pre-trained model achieves accuracies of 90.61%, 69.50% and 58.36% on the original CIFAR-10, CIFAR-100 and Tiny-ImageNet test sets respectively.
SNN-ResNet
To explore other CNN architecture, we use SNN-ResNet11 [27], which trains SNNs directly via surrogate gradient. The pre-trained model used in our experiments can achieve 90.95% accuracy on the original CIFAR-10 test sets.
Spike-driven Transformer V3
To evaluate the effectiveness of our work on more complex networks and datasets, we utilized a 19M parameter model in Spike-driven Transformer V3 [44], which achieves 79.60% accuracy on the ImageNet-1K validation set. This model has a patch size of 14 $\times$ 14, with an embedding dimension of 360 output by the final feature extractor. Instead of using the class token, the entire embedding feature map is employed as the input to the classifier. This design naturally aligns with our method, where the channel dimension $C$ corresponds to the embedding dimension, and the spatial dimension $D$ matches the patch size.
SNN-ConvLSTM
DVS Gesture is a neuromorphic dataset with a temporal dimension, making it well-suited for evaluating RNN-based architectures. To validate the effectiveness of our work on such networks, we employed a spike-driven 2ConvLSTM-1FC model, with hidden layer dimensions of 32 and 64, respectively. This network achieves an accuracy of 93.36% on the original DVS Gesture test set.
D.3 More Details on Implementation
To better reflect real-world scenarios, we adapt our method to independently perform adaptation on each individual test input. In the same experiment, the learning rate remains fixed across different shifted domains, as the domain of a given test sample is typically unknown in real-world settings. For each set of experiments, we employ the SGD optimizer and evaluate learning rates ranging from 0.001 to 0.8, selecting the optimal value based on performance. We trained and tested all models and datasets in 4 NVIDIA GeForce RTX 3090 GPUs.